# R2R: Efficiently Navigating Divergent Reasoning Paths with Small-Large Model Token Routing

### v1

@@ Abstract
Large Language Models (LLMs) achieve impressive reasoning capabilities at the cost of substantial inference overhead, posing substantial deployment challenges. Although distilled Small Language Models (SLMs) significantly enhance efficiency, their performance suffers as they fail to follow LLMs’ reasoning paths. Luckily, we reveal that only a small fraction of tokens genuinely diverge reasoning paths between LLMs and SLMs. Most generated tokens are either identical or exhibit neutral differences, such as minor variations in abbreviations or expressions. Leveraging this insight, we introduce Roads to Rome (R2R), a neural token routing method that selectively utilizes LLMs only for these critical, path-divergent tokens, while leaving the majority of token generation to the SLM. We also develop an automatic data generation pipeline that identifies divergent tokens and generates token-level routing labels to train the lightweight router. We apply R2R to combine R1-1.5B and R1-32B models from the DeepSeek family, and evaluate on challenging math, coding, and QA benchmarks. With an average activated parameter size of 5.6B, R2R surpasses the average accuracy of R1-7B by 1.6$\times$, outperforming even the R1-14B model. Compared to R1-32B, it delivers a 2.8$\times$ wall-clock speedup with comparable performance, advancing the Pareto frontier of test-time scaling efficiency. Our code is available at https://github.com/thu-nics/R2R.
@@ 1 Introduction
Large Language Models (LLMs) demonstrate strong capabilities across a wide range of tasks [1, 2, 3]. Building upon the largest and strongest LLMs, test-time scaling has become a prominent way to further boost their abilities on challenging tasks [4, 5, 6, 7]. It is typically done by generating extensive Chain-of-Thought (CoT) reasoning before producing the final answer. However, this approach requires massive LLMs with hundreds of parameters to generate thousands of tokens per query [8], resulting in significant inference overhead [9].
Distilled Small Language Models (SLMs), containing only a few billion parameters, offer much higher generation efficiency. Through supervised finetuning on LLM responses, SLMs can mimic LLM reasoning behaviors, making them a popular alternative. However, SLMs may still produce different reasoning paths from their LLM counterparts during inference, causing severe performance degradation. For example, compared to the R1-32B LLM, the R1-1.5B SLM provides different final answers on 45% questions in the AIME benchmark [10], suffering a 4.8$\times$ reduction in accuracy as shown in Table 2.
Figure 1: (a) Examples of R2R routing objective. Given a partial response as context, if SLM next-token prediction is not identical with LLM’s, it is further categorized as neutral or divergent based on their affects on the reasoning path. (b) Distribution of identical, neural and divergent labels in the R2R training set with 7.6M token labels.
@@
If addressed, we could unlock substantial efficiency advantage of SLMs for most generation steps, yet preserving the high-quality reasoning typical of LLMs. This can enable better test-time scaling by advancing the efficieny-performance Pareto frontier.
The main challenge of SLM-LLM mix inference involves two interconnected parts: labeling the preferred model under certain objective, and designing the routing scheme to enforce it during inference. Previous methods typically route at the query level, selecting either SLM or LLM for entire response to maximize human preference win-rate within a cost budget [11, 12]. However, these approaches rely on human annotations and complex router designs, whose data labeling and routing scheme are both too expensive for fine-grained, token-level routing. Alternatively, speculative decoding methods aim for identical outputs between SLM and LLM at the token level [13, 14, 15, 16]. They draft outputs with SLMs (or draft models) and periodically verify them with LLMs. While accurate, this strict verification leads to low acceptance rates. Additionally, mid-draft differences invalidate all subsequent tokens, severely restricting the accepted lengths as shown in Figure 2(b).
@@
  • Performance-Efficiency Pareto Frontier. R2R enables more efficient performance scaling at test time. Compared to query-level routing and distilled R1-14B, it delivers 1.1–1.5$\times$ higher AIME accuracy with 1.5–1.6$\times$ lower latency. R2R also provides a 2.8$\times$ wall-clock speedup over R1-32B LLM at similar accuracy, while raising R1-1.5B SLM accuracy by 4.6$\times$ with only 12.9% LLM usage.
@@ 2 Related Work
Test-time scaling improves LLM performance at the higher cost of inference, often through the explicit generation of the CoT reasoning paths [4, 6]. For more effective scaling, previous works optimize the length and width of reasoning paths, or reduce the generation overhead of each path [17, 18].
Controlling Reasoning Paths. Some approaches reduce LLM output lengths. They employ prompting [19], post-training [20], or heuristics [21] to generate concise CoT with fewer decoding steps [19, 21]. Others explore the width of paths. They let LLMs generate multiple reasoning paths in parallel, then select the best outcome with methods like best-of-N voting [22] or external verifiers [23]. Both strategies modify the structure of reasoning paths, which are perpendicular to R2R’s focus on reducing the overhead of each path.
Model Routing. Model routing reduces generation cost by selecting the most suitable model for each query based on difficulty and budget. Current works explore selection criteria of learned human preferences [11], reward signals [24], query tags [25], and model profiles [12]. Despite simplicity, they enforce the same LLM for each response, yielding suboptimal performance for the common mixed-difficulty generations. In contrast, R2R routes at token level to further improve efficiency.
Speculative Decoding. Speculative decoding accelerates generation by fixing the low parallelism in LLM decoding [13, 14, 16, 15]. It drafts outputs through SLM sequential decoding, then periodically verifies them with high-parallel LLM prefilling. However, speculative decoding pursuits identical output token (distribution) between SLM and LLM, causing low acceptance rate. In addition, it is often that not all tokens generated by SLM within one draft-verify cycle can pass LLM verification. The mid-draft rejection invalidates all subsequent drafts and LLM verifications as shown in Figure 2(b), leading to frequent rollbacks. Expanding the single draft-chain to draft-tree alleviative the problem, but also incurs higher overheads that harm batch serving efficiency [16]. Considering the internality of CoT process, R2R accepts neutrally different output tokens, and immediately corrects all divergent tokens to avoid any rollback.
@@ 3.2 Path-following Routing Strategy
$m_{i}$ =$=\begin{cases}\theta_{s},&\underbrace{y_{i}(\theta_{s}|S_{<i})=y_{i}(\theta_{l}|S_{<i})}_{\textit{identical}}\;\;\text{or}\;\;\underbrace{\mathcal{V}(\mathcal{S}_{s},\mathcal{S}_{l})=1}_{\textit{neutral}}\\ \theta_{l},&\underbrace{\mathcal{V}(\mathcal{S}_{s},\mathcal{S}_{l})=0}_{\textit{divergent}}\end{cases}$ (3)
@@
In practice, we use sentence-level path-following routing, where the continuation ends at the current sentence, as shown in Figure 3 (step 2). We monitor sentence-ending symbols, like period, during continuation and use existing semantical sentence seperators [26, 27] to conclude generation if the sentence truly ends. To verify this local continuation, a capable LLM serves as a sentence-level verifier $\mathcal{V}^{\prime}$, as shown in Figure 3 (step 3). It is prompted to compare the continuations and determine whether the initial token difference introduces a meaningful divergence from the LLM’s intended reasoning path, or merely a neutral abbreviation. Instead of verifying the entire generation, this approach checks the reasoning path at the sentence level, greatly improving data labeling efficiency.
We empirically validate the effectiveness of sentence-level path-following routing using Qwen2.5-72B [28] as the verifier model, with prompts detailed in Appendix F.1. Among 17 AIME-24 questions correctly solved by R1-32B within an 8K-token limit, our path-following strategy achieves comparable accuracy (16 questions correctly answered) while relying on the smaller R1-1.5B model for only 3% of generated tokens.
@@ 4.1 Training Data Generation
Figure 3 shows our data generation pipeline. Given queries from existing datasets, we first obtain the complete LLM response, either directly from the dataset or via batched LLM inference. Next, we use highly parallel SLM prefilling to efficiently identify tokens where the SLM prediction is identical to the LLM, allowing us to exclude about 90% of tokens from further processing. For the remaining 10% of differing tokens, we perform batched LLM continuations from each SLM prediction. To further improve efficiency, we apply prefix caching in current frameworks [29, 30] to reuse KV-Cache computations for shared context prefixes across multiple continuations (e.g., everything preceding Let in Figure 3). Continuations for the corresponding LLM tokens, $\mathcal{S}(\theta_{l})$, are directly extracted from the pre-generated LLM response, eliminating redundant computation. Finally, the verifier model compares both continuations and label routing preference.
Using this pipeline, we efficiently generate 7.6 million routing labels in approximately 2.3 days on 8 A800 GPUs, covering topics of math, coding, and QA with queries from the Bespoke-Stratos [31] dataset. Table 1 summarizes the statistics of the generated training dataset.
@@ 4.2 Predictive Indicators of Divergence
SLM Logits. As shown in Figure 4(a), divergent tokens exhibit substantially higher entropy in the SLM’s output logits, with a 3.8$\times$ mean value over that of non-divergent tokens. We observe similar trends with other uncertainty measures, such as aleatoric uncertainty [32]. These empirical results indicate that increased uncertainty in SLM predictions is strongly correlated with token divergence. Motivated by this, our router takes top-100 SLM logit values as one of its input features.
Token Frequency. Figure 4(b) shows that low-frequency tokens in the dataset are more likely to be divergent. This likely arises from the long-tail token distribution in the training data, making rare tokens harder for SLMs to model effectively due to the limited capacity [33]. Given this insight, our router explicitly incorporates token-frequency biases by using the token embedding of as router inputs.
@@ 4.3 Router Design and Routing Scheme
Model Architecture. Guided by insights in Section 4.2, we design the neural router as a lightweight, five-layer feed-forward network (FFN) with 56M parameters. It takes the SLM’s output logits and tokenized embedding, along with its last-layer hidden states for additional semantic context. All inputs are linearly projected, concatenated, and fed into the FFN backbone. The router outputs a binary classification probability, indicating whether the current token diverges from the LLM’s reasoning path. Full network architecture detail descriptions are in Appendix A.1.
@@
Routing Scheme. Unlike speculative decoding methods that periodically verify SLM outputs, our routing scheme aims to immediately decide whether to accept each SLM token, eliminating the need for rollbacks. As shown in Figure 2, this approach reduces unnecessary draft and verification computations, which is especially beneficial in computation-intensive batch-serving scenarios. Specifically, the neural router estimates divergence probability at each generation step using SLM outputs. When this probability exceeds a predefined threshold $p_{\text{th}}$, the LLM is invoked to correct the current output token. Following speculative decoding methods [13, 15], we utilize highly parallel prefilling for efficient LLM KV-Cache updates, whose overhead can be further reduced by overlapping them with SLM decoding [34].
@@ 5.1 Setup
Baselines. We use DeepSeek-R1-Distill-Qwen models as baselines, denoted by R1-$M$B, where $M$ indicates the model size in billions. We designate R1-1.5B and R1-32B as SLM and LLM, respectively, while intermediate sizes (7B, 14B) capture distillation scaling behavior. We compare various query-level routing (QR) methods from the RouteLLM framework [11], including similarity-weighted ranking (QR-SW), matrix factorization (QR-MF), a BERT-based classifier (QR-BERT), and a Llama3-8B-based classifier (QR-LLM). For speculative decoding, we adopt EAGLE2 [14] and HASS [15] with R1-32B LLM. We use the official HASS draft model, and train the EAGLE2 draft model using its official script, as no pre-trained EAGLE2 draft is provided for R1-32B.
@@
Benchmark. We evaluate methods across challenging reasoning benchmarks, including mathematics (AIME 2024–2025 [10]; denoted as AIME), graduate-level question-answering (GPQA-Diamond [35]; denoted as GPQA), and coding tasks (LiveCodeBench 2024-08–2025-01; denoted as LiveCodeBench [36]). All experiments use a maximum output length of 32K tokens and zero generation temperature to ensure reproducibility.
Efficiency metric. Following previous works [37, 2], we use the average activated parameters per token as a hardware-agnostic efficiency metric, referred to as the average parameter ($\bar{M}$) for brevity. For query-level routing, $\bar{M}$ is computed as the weighted average of SLM and LLM parameters based on their activation ratios across all outputs. For R2R, $\bar{M}$ includes the SLM and router parameters, along with the LLM parameters weighted by the LLM activation ratio. We also report the total Cost ($C$), defined as average activated parameters multiplied by the average output tokens per query. The average parameter and total cost reflect the average decoding speed and total latency, respectively. In addition, we report hardware-specific decoding speed on NVIDIA A800-80GB GPUs using SGLang [29] framework.
@@ 5.2 Performance
Numerical Comparison. Table 2 shows numerical details of model performance around average parameter sizes of 7B and 14B. With an average parameter size of 5.6B, R2R outperforms the best query-level routing methods (in both 7B and 14B) by 1.4–2.4$\times$ and 1.2–1.4$\times$, respectively. Compared to distilled models, R2R improves accuracy by 1.4–1.7$\times$ over R1-7B and even surpasses R1-14B in average accuracy by 1.1$\times$. Relative to the extremes, R2R achieves 4.6$\times$ higher accuracy than R1-1.5B and retains 92% of R1-32B’s accuracy, while using the LLM for only 11–15% of tokens.
5.3 Wall-clock Speed
@@
Table 3 reports the wall-clock latency and speed for all methods on the AIME benchmark. All baselines use the official, highly efficient SGLang [29] framework and are evaluated with tensor parallelism on two NVIDIA A800-80GB GPUs. R2R uses the same thresholds as in Table 2; query-level routing methods use the 14B version for comparable performance. R2R achieves 1.62$\times$ and 2.76$\times$ generation speed over R1-14B and R1-32B, respectively. Compared to query-level routing, R2R delivers 1.48–1.52$\times$ speedup. It also outperforms highly optimized speculative decoding methods with tree-like drafts, which speedup mostly at the current single-batch setup [16]. Further system-level optimization can be done to yield even greater gains for R2R. Note that, in theory, speculative decoding should exactly match the R1-32B LLM’s output. However, we occasionally observe inconsistencies for very long outputs, likely due to cumulative numerical errors. We faithfully report this observation without questioning the equivalence guarantee of speculative decoding.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Large Language Models (LLMs) demonstrate strong capabilities across a wide range of tasks [achiam2023gpt4, liu2024deepseekv3, team2023gemini]. Building upon the largest and strongest LLMs, test-time scaling has become a prominent way to further boost their abilities on challenging tasks [jaech2024openai-o1, openai2025openai-o3, guo2025deepseek-r1, qwen3]. It is typically done by generating extensive Chain-of-Thought (CoT) reasoning before producing the final answer. However, this approach requires massive LLMs with hundreds of billions of parameters to generate thousands of tokens per query [openr1], resulting in significant inference overhead [deepseek2025inference].
[…]
Figure 1: (a) Examples of R2R routing objective. Given a partial response as context, if SLM next-token prediction is not identical with LLM’s, it is further categorized as neutral or divergent based on their effects on the reasoning path. (b) Distribution of identical, neutral and divergent labels in the R2R training set with 7.6M token labels.
@@
If addressed, we could unlock substantial efficiency advantage of SLMs for most generation steps, yet preserving the high-quality reasoning typical of LLMs. This can enable better test-time scaling by advancing the efficiency-performance Pareto frontier.
The main challenge of SLM-LLM mix inference involves two interconnected parts: labeling the preferred model under certain objective, and designing the routing scheme to enforce it during inference. Previous methods typically route at the query level, selecting either SLM or LLM for entire response to maximize human preference win-rate within a cost budget [ong2024routellm, feng2025graphrouter]. However, these approaches rely on human annotations and complex router designs, whose data labeling and routing scheme are both too expensive for fine-grained, token-level routing. Alternatively, speculative decoding methods aim for identical outputs between SLM and LLM at the token level [li2024eagle, li2024eagle2, zhang2024hass, li2025eagle3]. They draft outputs with SLMs (or draft models) and periodically verify them with LLMs. While accurate, this strict verification leads to low acceptance rates. Additionally, mid-draft differences invalidate all subsequent tokens, severely restricting the accepted lengths as shown in Figure 2(b).
@@
[…]
@@ 2 Related Work
Test-time scaling improves LLM performance at the higher cost of inference, often through the explicit generation of the CoT reasoning paths [jaech2024openai-o1, guo2025deepseek-r1]. For more effective scaling, previous works optimize the length and width of reasoning paths, or reduce the generation overhead of each path [qu2025reasonsurvey, sui2025reasoningsurvey2].
Controlling reasoning paths. Some approaches reduce LLM output lengths. They employ prompting [xu2025cod], post-training [han2024tokenbudget], or heuristics [wang2025overthink] to generate concise CoT with fewer decoding steps [xu2025cod, wang2025overthink]. Others explore the width of paths. They let LLMs generate multiple reasoning paths in parallel, then select the best outcome with methods like best-of-N voting [lightman2023verify] or external verifiers [xin2024deepseekprover]. Both strategies modify the structure of reasoning paths, which are perpendicular to R2R’s focus on reducing the overhead of each path.
Model routing. Model routing reduces generation cost by selecting the most suitable model for each query based on difficulty and budget. Current works explore selection criteria of learned human preferences [ong2024routellm], reward signals [lu2023zooter], query tags [wang2025mixllm], and model profiles [feng2025graphrouter]. Despite simplicity, they enforce the same LLM for each response, yielding suboptimal performance for the common mixed-difficulty generations. In contrast, R2R routes at token level to further improve efficiency.
Speculative decoding. Speculative decoding accelerates generation by fixing the low parallelism in LLM decoding [li2024eagle, li2024eagle2, li2025eagle3, zhang2024hass]. It drafts outputs through SLM sequential decoding, then periodically verifies them with high-parallel LLM prefilling. However, speculative decoding pursues identical output token (distribution) between SLM and LLM, causing low acceptance rate. In addition, it is often that not all tokens generated by SLM within one draft-verify cycle can pass LLM verification. The mid-draft rejection invalidates all subsequent drafts and LLM verifications as shown in Figure 2(b), leading to frequent rollbacks. Expanding the single draft-chain to draft-tree alleviates the problem, but also incurs higher overheads that harm batch serving efficiency [li2025eagle3]. Considering the internality of CoT process, R2R accepts neutrally different output tokens, and immediately corrects all divergent tokens to avoid any rollback.
@@ 3.2 Path-following Routing Strategy
$m_{i}$ $=\begin{cases}\theta_{s},&\underbrace{y_{i}(\theta_{s}|S_{<i})=y_{i}(\theta_{l}|S_{<i})}_{\textit{identical}}\;\;\text{or}\;\;\underbrace{\mathcal{V}(\mathcal{S}_{s},\mathcal{S}_{l})=1}_{\textit{neutral}}\\ \theta_{l},&\underbrace{\mathcal{V}(\mathcal{S}_{s},\mathcal{S}_{l})=0}_{\textit{divergent}}\end{cases}$ (3)
@@
In practice, we use sentence-level path-following routing, where the continuation ends at the current sentence, as shown in Figure 3 (step 2). We monitor sentence-ending symbols, like period, during continuation and use existing semantical sentence separators [loper2002nltk, sun_jieba] to conclude generation if the sentence truly ends. To verify this local continuation, a capable LLM serves as a sentence-level verifier $\mathcal{V}^{\prime}$, as shown in Figure 3 (step 3). It is prompted to compare the continuations and determine whether the initial token difference introduces a meaningful divergence from the LLM’s intended reasoning path, or merely a neutral abbreviation. Instead of verifying the entire generation, this approach checks the reasoning path at the sentence level, greatly improving data labeling efficiency.
We empirically validate the effectiveness of sentence-level path-following routing using Qwen2.5-72B [yang2024qwen2] as the verifier model, with prompts detailed in Appendix F.1. Among 17 AIME-24 questions correctly solved by R1-32B within an 8K-token limit, our path-following strategy achieves comparable accuracy (16 questions correctly answered) while relying on the larger R1-32B model for only 3% of generated tokens.
@@ 4.1 Training Data Generation
Figure 3 shows our data generation pipeline. Given queries from existing datasets, we first obtain the complete LLM response, either directly from the dataset or via batched LLM inference. Next, we use highly parallel SLM prefilling to efficiently identify tokens where the SLM prediction is identical to the LLM, allowing us to exclude about 90% of tokens from further processing. For the remaining 10% of differing tokens, we perform batched LLM continuations from each SLM prediction. To further improve efficiency, we apply prefix caching in current frameworks [zheng2023sglang, kwon2023vllm] to reuse KV-Cache computations for shared context prefixes across multiple continuations (e.g., everything preceding Let in Figure 3). Continuations for the corresponding LLM tokens, $\mathcal{S}(\theta_{l})$, are directly extracted from the pre-generated LLM response, eliminating redundant computation. Finally, the verifier model compares both continuations and label routing preference. Further analysis of the robustness of the verifier and the comparison to human experts are provided in Appendix B.1.
Using this pipeline, we efficiently generate 7.6 million routing labels in approximately 2.3 days on 8 A800 GPUs, covering topics of math, coding, and QA with queries from the Bespoke-Stratos [bespokestratos] dataset. Table 1 summarizes the statistics of the generated training dataset.
@@ 4.2 Predictive Indicators of Divergence
[…]
Token frequency. Figure 4(b) shows that low-frequency tokens in the dataset are more likely to be divergent. This likely arises from the long-tail token distribution in the training data, making rare tokens harder for SLMs to model effectively due to the limited capacity [zevallos2023frequency]. Given this insight, our router explicitly incorporates token-frequency biases by using the token embedding of as router inputs.
@@ 4.3 Router Design and Routing Scheme
Model architecture. Guided by insights in Section 4.2, we design the neural router as a lightweight, six-layer feed-forward network (FFN) with 56M parameters. It takes the SLM’s output logits and tokenized embedding, along with its last-layer hidden states for additional semantic context. All inputs are linearly projected, concatenated, and fed into the FFN backbone. The router outputs a binary classification probability, indicating whether the current token diverges from the LLM’s reasoning path. Full network architecture detail descriptions are in Appendix A.1.
@@
Routing scheme. Unlike speculative decoding methods that periodically verify SLM outputs, our routing scheme aims to immediately decide whether to accept each SLM token, eliminating the need for rollbacks. As shown in Figure 2, this approach reduces unnecessary draft and verification computations, which is especially beneficial in computation-intensive batch-serving scenarios. Specifically, the neural router estimates divergence probability at each generation step using SLM outputs. When this probability exceeds a predefined threshold $p_{\text{th}}$, the LLM is invoked to correct the current output token. Following speculative decoding methods [li2024eagle, zhang2024hass], we utilize highly parallel prefilling for efficient LLM KV-Cache updates, whose overhead can be further reduced by overlapping them with SLM decoding [akhauri2025splitreason].
@@ 5.1 Setup
Baselines. We use DeepSeek-R1-Distill-Qwen models as baselines, denoted by R1-MMB, where $M$ indicates the model size in billions. We designate R1-1.5B and R1-32B as SLM and LLM, respectively, while intermediate sizes (7B, 14B) capture distillation scaling behavior. We compare various query-level routing (QR) methods from the RouteLLM framework [ong2024routellm], including similarity-weighted ranking (QR-SW), matrix factorization (QR-MF), a BERT-based classifier (QR-BERT), and a Llama3-8B-based classifier (QR-LLM). For speculative decoding, we adopt EAGLE2 [li2024eagle2] and HASS [zhang2024hass] with R1-32B LLM. We use the official HASS draft model, and train the EAGLE2 draft model using its official script, as no pre-trained EAGLE2 draft is provided for R1-32B.
@@
Benchmark. We evaluate methods across challenging reasoning benchmarks, including mathematics (AIME 2024–2025 [aimedata]; denoted as AIME), graduate-level question-answering (GPQA-Diamond [rein2023gpqa]; denoted as GPQA), and coding tasks (LiveCodeBench 2024-08–2025-01; denoted as LiveCodeBench [jain2024livecodebench]). All experiments use a maximum output length of 32K tokens and zero generation temperature to ensure reproducibility.
Efficiency metric. Following previous works [jiang2024mixtral, liu2024deepseekv3], we use the average activated parameters per token as a hardware-agnostic efficiency metric, referred to as the average parameter ($\bar{M}$) for brevity. For query-level routing, $\bar{M}$ is computed as the weighted average of SLM and LLM parameters based on their activation ratios across all outputs. For R2R, $\bar{M}$ includes the SLM and router parameters, along with the LLM parameters weighted by the LLM activation ratio. We also report the total Cost ($C$), defined as average activated parameters multiplied by the average output tokens per query. The average parameter and total cost reflect the average decoding speed and total latency, respectively. In addition, we report hardware-specific decoding speed on NVIDIA A800-80GB GPUs using SGLang [zheng2023sglang] framework.
@@ 5.2 Performance
[…]
Generalizability. Beyond the R1 family, we further train and evaluate R2R on the Qwen3 series (both dense and MoE variants) spanning 0.6B, 1.7B, 8B, 30B-A3B, and 32B parameter scales. Across all configurations, R2R consistently surpasses both the base Qwen3 models and query-level routing baselines, demonstrating strong adaptability and cross-architecture generalization. We also evaluate on Arena-Hard (dialogue) and MMLU-Redux-Philosophy benchmarks, which are beyond the selected mathematical reasoning, QA, and code generation tasks. Remarkably, R2R continues to outperform R1-14B while maintaining an average activated parameter size of only 6.1–6.7B, confirming its robust generalization across domains and model families (see Appendix  B.2.2).
Sampling method extension. Beyond greedy decoding, we extend R2R to nucleus (top-p = 0.95) with temperature (temperature = 0.6). R2R still greatly advances Pareto frontier, with 1.4-1.5x higher AIME score over query-level routing, reaching R1-14B score with only 8.6B average parameters. Implementation details and theoretical analysis are provided in Appendix C.1.
5.3 Efficiency
@@
[…]
Computation and memory access. Compared with the LLM baseline, R2R achieves a $5.4\times$ reduction in per-token memory access. The total computation increases only marginally due to the additional SLM and router computation. Because decoding is largely memory-bound, the reduced memory traffic yields higher throughput. This substantial reduction in memory traffic translates into higher throughput that outweighs the minor compute overhead. Compared with speculative decoding approaches such as Eagle2 and HASS, R2R requires about $17.0\times$ less total computation, primarily because its immediate correction prevents the LLM from repeatedly prefilling unused tokens during periodic verification (see Figure 2). At the same time, it reduces memory access by $2.4\text{–}2.5\times$ relative to these speculative methods, yielding a more balanced compute–memory trade-off. More details are provided in Appendix B.3.2.